# BE Lyncis: An Extremely Eccentric Binary with the Nearest Known Black Hole

### v1

@@ Abstract
We report the discovery of an exceptionally eccentric binary system, BE Lyncis (BE Lyn), which hosts the nearest known black hole (BH) to Earth. Through the analysis of TESS photometry combined with an extensive set of times of maximum light spanning 39 years, we identify BE Lyn as a high-amplitude $\delta$ Scuti star in a binary with an orbital period of $\approx 15.9$ years and an extraordinary orbital eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence)—the highest reliably measured for any binary system. Dynamical constraints impose an upper limit on the orbital inclination of $i\lesssim 4.0^{\circ}$, corresponding to a companion mass of $M_{2}\gtrsim 17.5~M_{\odot}$, which unequivocally favors a black hole. This system provides a unique laboratory for studying asteroseismology in strong gravitational fields, the formation of black holes via asymmetric supernovae, and the evolution of extreme binary systems. Our work demonstrates, for the first time, the successful application of the light-travel time effect in a pulsating variable to unveil a dormant black hole, establishing a novel method for BH detection in non-interacting binaries.
@@ I Introduction
Black holes (BHs) of stellar origin represent one of the most enigmatic predictions of general relativity. While several dozen stellar-mass BHs have been identified in X-ray binaries through accretion signatures (R. W. Romani, 1998), the population of dormant (non-accreting) BHs in detached binaries remains largely unexplored. Detecting such systems is challenging, as they typically reveal themselves only through subtle gravitational interactions with luminous companions. Recent astrometric surveys, such as Gaia, have uncovered a handful of such quiescent BH binaries (K. El-Badry et al., 2023; A. Tanikawa et al., 2023; Gaia Collaboration et al., 2024), yet they remain exceedingly rare. An independent, complementary detection method—leveraging the precise clocks provided by pulsating stars—offers a promising pathway to expand this census and probe different regions of parameter space.
@@
Pulsating $\delta$ Scuti stars are particularly well-suited for this purpose. These stars, with spectral types A–F, occupy the intersection of the main sequence and the classical Cepheid instability strip in the Hertzsprung–Russell diagram. Their oscillations, driven by the $\kappa$ mechanism operating in the partial ionization zone of helium, provide remarkably stable periodic signals with periods ranging from 15 minutes to 8 hours (T. Kallinger et al., 2008; G. Handler, 2009; D. B. Guenther et al., 2009; K. Uytterhoeven et al., 2011; D. L. Holdsworth et al., 2014; T. Steindl et al., 2022). The subclass of high-amplitude $\delta$ Scuti stars (HADS) exhibits photometric variations exceeding 0.1 mag and typically rotates slowly ($v\sin i\leq 30$ km s-1), making them ideal precision tools for detecting the minite period modulations induced by orbital motion in binary systems—the light-travel time effect (LTTE).
@@
BE Lyncis (BE Lyn; HD 79889) is a well-studied HADS with a mean magnitude of $\langle V\rangle=8.8$ mag, spectral type A3, and sub-solar metallicity ([Fe/H] = -0.598) (C. Kim et al., 2012; Abdurro’uf et al., 2022). It was first identified as a variable star by T. Oja (1986) and later confirmed as a HADS with a pulsation period of $\sim$0.09587 days (T. Oja, 1987). Subsequent studies of its period variations revealed anomalies in OCO-C diagrams, occasionally suggestive of a parabolic trend indicative of a binary companion (e.g., E. Rodriguez et al., 1990; Y. Liu et al., 1991; E. Wunder et al., 1992; C. Boonyarak et al., 2011). Several investigations indeed proposed the existence of a low-mass companion (L. L. Kiss & K. Szatmary, 1995; A. Derekas et al., 2009; C. Boonyarak et al., 2011; J. H. Peña et al., 2015; A. Liakos & P. Niarchos, 2017), though inferred orbital parameters varied significantly due to limitations in the temporal coverage and precision of the data.
@@
Two key developments now enable a definitive reassessment of this system. First, high-precision, continuous photometry from the Transiting Exoplanet Survey Satellite (TESS) allows for an exquisite characterization of the primary star’s pulsation modes and yields significantly more accurate times of maximum light. Second, modern computational techniques, such as robust multi-parameter global fitting via Markov Chain Monte Carlo (MCMC) methods, permit a physically consistent and precise determination of binary orbital parameters.
@@
Here, we present a comprehensive analysis of BE Lyn, combining TESS photometry with a newly compiled, 39-year baseline of times of maximum light. Our work not only confirms its binary nature but reveals an astrophysical gem: an ultra-wide, extremely eccentric orbit harboring a massive black hole companion. This system holds multiple distinctions: (i) it possesses the highest orbital eccentricity ever reliably measured in a binary; (ii) it hosts the closest known black hole to Earth, at a distance of approximately 250 pc (C. A. L. Bailer-Jones et al., 2021); (iii) it is the only known binary comprising a pulsating star and a quiescent black hole, opening a new window for asteroseismology in strong gravity; and (iv) its extreme orbital configuration places stringent constraints on the natal kick imparted during the black hole’s formation, challenging models of asymmetric supernova explosions. The detection methodology itself—using pulsation timing in a $\delta$ Scuti star—pioneers a new, independent channel for discovering dormant black holes.
@@ II.1 Pulsation Properties and the Stellar Clock
Analysis of the TESS photometry (Figure A1) reveals BE Lyn as a multi-periodic pulsator. We identify 37 significant frequencies, comprising 7 independent modes and 30 harmonics or combination frequencies (Figure A2, Table A1). The two dominant frequencies, $f_{0}=10.4309\pm 0.0005$ c/d and $f_{1}=13.4226\pm 0.0005$ c/d, exhibit a period ratio of $P_{1}/P_{0}=0.777$, unambiguously identifying them as the fundamental and first-overtone radial modes, respectively. The fundamental mode amplitude dominates the light curve, exceeding that of the first overtone by a factor of $\sim 45$ and the remaining independent modes by factors >700>700 (Figure A2). This clear amplitude hierarchy renders the $f_{0}$ pulsation a remarkably clean and stable clock, enabling the application of the classical OCO-C method to detect minute orbital perturbations with negligible contamination from multi-periodic effects. The five additional independent frequencies ($f_{2}$$f_{6}$), residing in the 9179-17 c/d range with low amplitudes, are likely non-radial p-modes. This rich pulsation spectrum establishes BE Lyn as a promising target for future detailed asteroseismic modeling.
@@ II.2 Orbital Solution: An Extraordinarily Eccentric Binary
The OCO-C analysis, incorporating a 39-year baseline of 445 times of maximum light, reveals a pronounced, long-period modulation superimposed on a secular period increase (Figure 1). Global fitting with a model combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term yields the orbital solution presented in Table 1.
@@
The most striking result is the extraordinary orbital eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence). With this value, BE Lyn claims the record for the highest reliably measured eccentricity in any binary system, surpassing previous benchmarks such as WDS 05354-3316 ($e=0.985\pm 0.002$ (A. Tokovinin et al., 2020)) and $\zeta$ Boo ($e=0.980450\pm 0.000064$ I. Waisberg et al. 2024, 2025). The orbital period is $P_{\mathrm{orb}}=5811^{+29}_{-30}$ days ($\approx 15.9$ years). The associated mass function, $f(m)=0.005^{+0.046}_{-0.004}M_{\odot}$, and small projected semi-major axis ($a\sin i=1.1^{+1.3}_{-0.4}$ AU) are indicative of a massive, unseen companion. This discovery underscores the power of long-term, precise photometric monitoring of pulsating stars to unveil extreme binary architectures.
II.3 Characterization of the Pulsating Primary
To place dynamical constraints in context, we determined the properties of the primary star through asteroseismic modeling. We computed a grid of stellar evolutionary tracks using MESA (Modules for Experiments in Stellar Astrophysics; B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023) and corresponding adiabatic pulsation frequencies using GYRE (R. H. D. Townsend & S. A. Teitler, 2013; R. H. D. Townsend et al., 2018; J. Goldstein & R. H. D. Townsend, 2020; M. Sun et al., 2023). The optimal model, selected by matching the observed $f_{0}$, $f_{1}$, and the intrinsic period variation rate $(1/P_{0})(dP_{0}/dt)$, is summarized in Table A2 and located on the Hertzsprung-Russell diagram in Figure 2.
The best-fit model places the primary on the post-main sequence with a mass of $M_{1}=1.55M_{\odot}$, effective temperature $\log T_{\mathrm{eff}}=3.902$, and luminosity $\log(L/L_{\odot})=1.326$. The agreement between the observed $(1.02\pm 0.08)\times 10^{-8}$ yr-1 and modeled ($1.34\times 10^{-8}$ yr-1) period variation rates is noteworthy. Unlike some HADS in binaries that exhibit anomalously large period variations due to mass transfer (e.g., H.-F. Xue & J.-S. Niu, 2020; M. Lian et al., 2025), BE Lyn’s primary appears consistent with single-star evolution. This suggests the system is currently detached and has not undergone significant mass exchange.
@@ II.3 Characterization of the Pulsating Primary
II.4 A Nearby Black Hole in an Extreme Orbit
Combining the orbital parameters with the primary mass ($M_{1}=1.55M_{\odot}$) introduces critical constraints from orbital stability: (i) the periastron separation must exceed the primary’s radius to avoid catastrophic collision; (ii) the primary’s Roche lobe radius must exceed the primary’s radius to avoid Roche-lobe overflow near the periastron. Applying these conditions, along with the mass function (Eq. A3), tightly constrains the system geometry (Figure 3). We derive an upper limit on the orbital inclination of $i\lesssim 4.0^{\circ}$, implying a lower limit on the companion mass of $M_{2}\gtrsim 17.5M_{\odot}$.
Figure 3: Dynamical constraints on the unseen companion. Top: Companion mass $M_{2}$ versus orbital inclination $i$, derived from the mass function (red curve). Bottom: Primary-companion separation at periastron versus $i$ (green curve). The intersection of the primary’s Roche lobe radius $R_{1,\mathrm{R}}$ (blue curve) and its physical radius $R_{1}$ (grey curve) sets the stringent upper limit $i\lesssim 4.0^{\circ}$ (dark blue dot), corresponding to the lower mass limit $M_{2}\gtrsim 17.5M_{\odot}$ (dark red dot). The gray hatched region denotes parameter space excluded.
A compact object of this mass cannot be supported by electron or neutron degeneracy pressure, ruling out a white dwarf or neutron star. The companion must therefore be a black hole. With a Gaia DR3 distance of approximately 250 pc (C. A. L. Bailer-Jones et al., 2021), this black hole becomes the nearest known to Earth, residing about half the distance of the previous record-holders Gaia BH1 (480 pc (K. El-Badry et al., 2023)) and Gaia BH3 (590 pc (Gaia Collaboration et al., 2024)). Its proximity offers unparalleled opportunities for detailed follow-up study.
The system’s most enigmatic feature—its extreme eccentricity—holds the key to its formation history. The absence of evidence for a dense stellar environment makes a dynamical scattering origin unlikely. Instead, the $e\approx 0.9989$ orbit is a natural signature of a strong natal kick imparted to the black hole during an asymmetric supernova explosion (H.-T. Janka, 2013; S. Popov et al., 2025). BE Lyn thus provides a pristine laboratory to constrain kick velocities and the physics of mass ejection in fallback supernovae, challenging models to account for the most eccentric black-hole binary known.
@@ II.5 Broader Implications and Future Prospects
BE Lyn represents a system of singular importance across multiple astrophysical domains. First, it is the only known binary pairing a pulsating star with a quiescent black hole. This unique configuration opens the novel field of strong-gravity asteroseismology, where the pulsations of the primary could, in principle, probe the local spacetime metric perturbed by the black hole companion, especially during the brief, close periastron passages.
Second, the discovery validates the pulsation-timing method as a powerful, independent technique for detecting dormant stellar-mass black holes in wide binaries. This approach is complementary to radial-velocity and astrometric surveys, being particularly sensitive to long-period systems where other methods require extended baselines.
Key questions remain. Precise dynamical mass measurements via future astrometric missions (e.g., Gaia, Theia) or very-long-baseline interferometry could pin down $M_{2}$ and $i$. Multi-wavelength searches for accretion signatures or bow-shock emission are warranted, despite the wide separation. Theoretical work is urgently needed to simulate binary evolution pathways that yield such extreme eccentricities post-supernova.
In conclusion, BE Lyn is not merely another binary system; it is a benchmark object. It challenges our understanding of black hole formation kicks, offers a new tool for black hole discovery, and provides a unique cosmic laboratory at our galactic doorstep.

### v2

# BE Lyncis: A Pulsating Star in the Most Eccentric Binary with a Massive Unseen Companion

@@ Abstract
We report the discovery of an exceptionally eccentric binary system, BE Lyncis (BE Lyn), which hosts a compact companion with mass $\gtrsim 2.5~M_{\odot}$. By combining TESS photometry with an extensive set of times of maximum light spanning 39 years, we identify BE Lyn as a high-amplitude $\delta$ Scuti star in a binary with an orbital period of $\approx 15.9$ years and an extraordinary eccentricity of $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence)—the most extreme eccentricity reliably measured for any binary system. Dynamical constraints limit the orbital inclination to $i\lesssim 10.1^{\circ}$, implying a companion mass $M_{2}\gtrsim 2.5~M_{\odot}$, which spans the mass gap between neutron stars and black holes. If the companion is a neutron star, it would be the most massive known; if a black hole, it would be the closest to Earth. This system provides a unique laboratory for studying asteroseismology in strong gravitational fields, as well as the formation and evolution of extremely eccentric binaries. Our work demonstrates the use of the light-travel time effect in a pulsating star to reveal a compact companion, offering a novel method for detecting black holes in non-interacting binaries.
software: astropy (Astropy Collaboration et al., 2022), Lightkurve (Lightkurve Collaboration et al., 2018), NumPy (C. R. Harris et al., 2020), SciPy (P. Virtanen et al., 2020), matplotlib (J. D. Hunter, 2007), emcee (D. Foreman-Mackey et al., 2013), MESA (B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023), GYRE (R. H. D. Townsend & S. A. Teitler, 2013)
@@ I Introduction
Pulsating $\delta$ Scuti stars, with spectral types A–F, occupy the intersection of the main sequence and the classical Cepheid instability strip in the Hertzsprung–Russell diagram. Their oscillations, driven by the $\kappa$ mechanism operating in the helium partial ionization zone, provide stable periodic signals with periods ranging from 15 minutes to 8 hours (T. Kallinger et al., 2008; G. Handler, 2009; D. B. Guenther et al., 2009; K. Uytterhoeven et al., 2011; D. L. Holdsworth et al., 2014; T. Steindl et al., 2022). High-amplitude $\delta$ Scuti stars (HADS), a subclass exhibiting photometric variations exceeding 0.1 mag and typically slow rotation ($v\sin i\leq 30$ km s-1), serve as precision celestial clocks. Their stability makes them ideally suited for detecting the minute period modulations induced by orbital motion in binary systems—a phenomenon known as the light-travel time effect (LTTE).
@@
BE Lyncis (BE Lyn; HD 79889) is a well-studied HADS with a mean magnitude of $\langle V\rangle=8.8$ mag, spectral type A3, and sub-solar metallicity ([Fe/H] = 0.598-0.598) (C. Kim et al., 2012; Abdurro’uf et al., 2022). First identified as a variable star by T. Oja (1986) and later confirmed as a HADS with a pulsation period of $\sim 0.09587$ days (T. Oja, 1987), it has been the subject of numerous period variation studies. Anomalies in its OCO-C diagram have long hinted at the presence of a binary companion, occasionally manifesting as a parabolic trend (e.g., E. Rodriguez et al., 1990; Y. Liu et al., 1991; E. Wunder et al., 1992; C. Boonyarak et al., 2011). Several investigations indeed proposed the existence of a low-mass companion (L. L. Kiss & K. Szatmary, 1995; A. Derekas et al., 2009; C. Boonyarak et al., 2011; J. H. Peña et al., 2015; A. Liakos & P. Niarchos, 2017), though the inferred orbital parameters varied substantially owing to limited temporal coverage and data precision.
@@
Two key developments now enable a definitive reassessment of this system. First, high-precision, continuous photometry from the Transiting Exoplanet Survey Satellite (TESS) allows exquisite characterization of the primary star’s pulsation modes and yields significantly more accurate times of maximum light. Second, modern computational techniques, such as robust multi-parameter global fitting via Markov Chain Monte Carlo (MCMC) methods, permit a physically consistent and precise determination of binary orbital parameters.
@@
Here we present a comprehensive analysis of BE Lyn, combining TESS photometry with a newly compiled, 39-year baseline of times of maximum light. Our work not only confirms its binary nature but reveals an astrophysical gem: an ultra-wide, extremely eccentric orbit harbouring a compact companion of mass $\gtrsim 2.5~M_{\odot}$. This places the companion spanning the mass gap between neutron stars and black holes. If it is a neutron star, it would be the most massive known; if a black hole, it would be the closest to Earth, at a distance of approximately 250 pc (C. A. L. Bailer-Jones et al., 2021).
The system holds multiple distinctions: (i) it possesses the highest orbital eccentricity ever reliably measured in any binary; (ii) it is the only known binary comprising a pulsating star and a massive compact companion, opening a new window for asteroseismology in strong gravitational fields; (iii) its extreme eccentricity ($e=0.9989$) provides a stringent observational test for models of supernova natal kicks and post-supernova mass transfer, offering critical insights into the formation pathways of such extreme binaries; and (iv) it validates the LTTE method using pulsating stars as a novel, independent pathway for discovering quiescent compact objects in non-interacting binaries. BE Lyn thus stands as a benchmark system, with profound implications for stellar evolution, black hole demographics, and strong-gravity astrophysics.
@@ II.1 Pulsation Properties and the Stellar Clock
Analysis of the TESS photometry reveals BE Lyn as a multi-periodic pulsator. We identify 32 significant frequencies, comprising 7 independent modes and 25 harmonics or combination frequencies (see Appendix A for more details). The two dominant frequencies, $f_{0}=10.4309\pm 0.0005$ c/d and $f_{1}=13.4226\pm 0.0005$ c/d, exhibit a period ratio $P_{1}/P_{0}=0.777$, unambiguously identifying them as the fundamental and first-overtone radial modes, respectively (W. Xue et al., 2023). The amplitude of the fundamental mode dominates the light curve, exceeding that of the first overtone by a factor of $\sim 45$ and those of the remaining independent modes by factors >700>700. This clear amplitude hierarchy makes the $f_{0}$ pulsation an exceptionally clean and stable clock, enabling the classical OCO-C method to detect minute orbital perturbations with negligible contamination from multi-periodic effects. The five additional independent frequencies ($f_{2}$$f_{6}$), which lie in the 9179-17 c/d range with low amplitudes, are likely non-radial p-modes. This rich pulsation spectrum establishes BE Lyn as a promising target for future detailed asteroseismic modeling.
@@ II.2 Orbital Solution: An Extraordinarily Eccentric Binary
The OCO-C analysis, incorporating a 39-year baseline of 445 times of maximum light, reveals a pronounced long-period modulation superimposed on a secular period increase (Figure 1). Global fitting with a model combining a quadratic term (intrinsic period variation) and a light-travel time effect (LTTE) term yields the orbital solution presented in Table 1.
@@
The most striking result is the extraordinary orbital eccentricity: $e=0.9989^{+0.0008}_{-0.0021}$ (>0.9968>0.9968 at 95% confidence). With this value, BE Lyn sets a new record for the highest reliably measured eccentricity in any binary system, surpassing previous benchmarks such as WDS 05354-3316 ($e=0.985\pm 0.002$ (A. Tokovinin et al., 2020)) and $\zeta$ Boo ($e=0.980450\pm 0.000064$ (I. Waisberg et al., 2024, 2025)). The orbital period is $P_{\mathrm{orb}}=5811^{+29}_{-30}$ days ($\approx 15.9$ years). The associated mass function, $f(m)=0.005^{+0.046}_{-0.004}M_{\odot}$, and small projected semi-major axis ($a\sin i=1.1^{+1.3}_{-0.4}$ AU) point to a massive, unseen companion. This discovery highlights the power of long-term, precise photometric monitoring of pulsating stars to uncover extreme binary architectures.
II.3 Asteroseismic Properties of the Primary Star
To place dynamical constraints in context, we determined the properties of the primary star through asteroseismic modeling. We computed a grid of stellar evolutionary tracks using MESA (Modules for Experiments in Stellar Astrophysics; B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023) and corresponding adiabatic pulsation frequencies using GYRE (R. H. D. Townsend & S. A. Teitler, 2013; R. H. D. Townsend et al., 2018; J. Goldstein & R. H. D. Townsend, 2020; M. Sun et al., 2023). The optimal model, selected by matching the observed frequencies $f_{0}$ and $f_{1}$, is summarized in Table 3 and located on the Hertzsprung-Russell diagram in Figure 2.
@@ II.3 Asteroseismic Properties of the Primary Star
The best-fit model places the primary on the post-main sequence, with a mass of $M_{1}=1.61M_{\odot}$, effective temperature $\log T_{\mathrm{eff}}=3.915$, and luminosity $\log(L/L_{\odot})=1.391$. Notably, the observed period variation rate, $(1.02\pm 0.08)\times 10^{-8}$ yr-1, differs from the model-predicted rate of $8.36\times 10^{-8}$ yr-1. This discrepancy may indicate ongoing mass transfer during the observed time interval.
II.4 A Massive Compact Companion in an Extreme Orbit
Combining the orbital parameters with the primary mass ($M_{1}=1.61M_{\odot}$) introduces a critical constraint from orbital stability: the periastron separation must exceed the primary’s radius to avoid catastrophic collision. This condition tightly constrains the system geometry (Figure 3). We derive an upper limit on the orbital inclination of $i\lesssim 10.1^{\circ}$, implying a lower limit on the companion mass of $M_{2}\gtrsim 2.5M_{\odot}$.
Figure 3: Dynamical constraints on the unseen companion. Top: Companion mass $M_{2}$ versus orbital inclination $i$, derived from the mass function (red curve). Bottom: Primary-companion separation at periastron versus $i$ (green curve). The intersection of the primary’s radius $R_{1}$ (grey curve) and primary-companion separation (green curve) sets the stringent upper limit $i\lesssim 10.1^{\circ}$ (dark green dot), corresponding to the lower mass limit $M_{2}\gtrsim 2.5M_{\odot}$ (dark red dot). The gray hatched region denotes parameter space excluded.
While natal kicks during an asymmetric supernova explosion can impart a relatively high eccentricity to binary systems (H.-T. Janka, 2013; S. Popov et al., 2025), the extreme eccentricity and semi-major axis of this system may require additional mechanisms to explain its later evolution. Interestingly, at periastron, the primary’s radius inevitably exceeds its Roche lobe radius (using the approximation of P. P. Eggleton (1983); see Figure 3), indicating that the primary must fill its Roche lobe, leading to mass transfer near periastron. [^1] Here, we thank P. Nagarajan et al. (2026) for pointing out an error in our previous calculation of the primary’s Roche lobe radius. This periastron mass transfer may have driven the elongation of the semi-major axis and the extreme eccentricity (A. Parkosidis et al., 2026, 2025), consistent with their unified semi-analytical framework and representing an extreme case of such evolution.
A compact object with mass $M_{2}\gtrsim 2.5M_{\odot}$ cannot be supported by electron degeneracy pressure, ruling out a white dwarf. Its mass places it squarely spanning the mass gap between neutron stars and black holes.
If the companion is a neutron star, it would challenge the current record for the most massive neutron star (PSR J0952-0607 at $2.35\pm 0.17M_{\odot}$ (R. W. Romani et al., 2022)). If instead it is a black hole, it would become the nearest known black hole to Earth, at a Gaia DR3 distance of approximately 250 pc (C. A. L. Bailer-Jones et al., 2021)—roughly half the distance of the previous record-holders Gaia BH1 (480 pc (K. El-Badry et al., 2023)) and Gaia BH3 (590 pc (Gaia Collaboration et al., 2024)). Its proximity offers unparalleled opportunities for detailed follow-up observations, including astrometric monitoring and multi-wavelength searches for accretion signatures.
@@ II.5 Broader Implications and Future Prospects
BE Lyn is a system of singular importance across multiple astrophysical frontiers. First, it is the only known binary comprising a pulsating star and a quiescent compact object spanning the mass gap. This unique configuration opens the door to strong-gravity asteroseismology: in principle, the primary’s pulsations could probe the local spacetime metric perturbed by the massive companion, especially during the brief, close periastron passages.
Second, this discovery validates the pulsation-timing method as a powerful and independent technique for detecting dormant stellar-mass compact objects in wide binaries. Complementary to radial-velocity and astrometric surveys, this approach is particularly sensitive to long-period systems where other methods require extended baselines.
Several key questions remain. Precise dynamical mass measurements via future astrometric missions (e.g., Gaia, Theia) or very-long-baseline interferometry could determine $M_{2}$ and $i$ unambiguously. Multi-wavelength searches for accretion signatures or bow-shock emission are warranted, despite the wide separation. Theoretical work is urgently needed to model binary evolution pathways that produce such extreme eccentricities post-supernova.
In conclusion, BE Lyn is not merely another binary system; it is a benchmark object. It challenges our understanding of black hole formation kicks, offers a new tool for discovering compact objects, and provides a unique cosmic laboratory at our galactic doorstep.